# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Next, we simulate a state distillation factory. In order to perform a classical simulation of a full factory, we focus on distillation of the $|{\overline{Y}}\rangle=\overline{S}|{\overline{+}}\rangle$ state (Fig. 4(a)), which allows the easy implementation of $\overline{S}$ gates in the surface code. Since this circuit has a similar structure to the practically relevant $|{\overline{T}}\rangle$ magic state distillation factories (Methods, ED Fig. 3), we expect them to have similar performance. We fix the error rate of the circuit to $p=0.1\%$, and vary the input infidelity $P_{\mathrm{in}}$ in Fig. 4(c). Examining the output $|{\overline{Y}}\rangle$ of a one-level factory, we find that as the code distance is increased, the output logical error rate $P_{\mathrm{out}}$ approaches the fidelity expected for ideal Clifford logical gates in the factory $P_{\mathrm{out}}=7P_{\mathrm{in}}^{3}+O(P_{\mathrm{in}}^{4})$ (see Methods for the full expression), across the explored fidelity regime.
@@
Finally, we simulate the logical error rate for a two-level $|{\overline{Y}}\rangle$ state distillation factory, involving a total of 113 logical qubits, where the output $|{\overline{Y}}\rangle$ states of a $d_{1}=5$ factory is fed into a second factory with $d_{2}=9$, with the distance chosen such that the logical error is dominated by the input state infidelity. As shown in Fig. 4(d), the logical error rates at each level of the distillation procedure are consistent with that expected based on the ideal factory formula (Methods), confirming that our approach is FT. Since the state injection procedure is agnostic to the particular state that is injected, we expect that our results will readily generalize to the setting of $|{\overline{T}}\rangle$ magic state factories.
@@
Transversal operations and correlated decoding were recently found to be highly effective in experiments with reconfigurable neutral atom arrays [36]. The principles of algorithmic fault-tolerance described here are the core underlying mechanisms of these observations, such as correlated decoding of a logical Bell state [36], and our results here indicate that the same techniques allow for $\Theta(d)$ time reduction for universal computation. While recent work has provided strong evidence that this reduction might be possible for circuits consisting purely of Clifford gates and Pauli basis inputs [15], up to now it has generally been believed that this conclusion does not hold when performing universal quantum computation [30, 42, 43], which crucially relies on the use of magic states and feed-forward operations. The present work not only demonstrates that this $\Theta(d)$ time cost reduction is broadly applicable to universal quantum computing, but also provides a theoretical foundation for it through mathematical fault tolerance proofs.
@@
Although our analysis focused on the use of an MLE decoder, our numerical simulations suggest that algorithms with polynomial runtime can still achieve a competitive threshold [41], and the development of improved, parallel correlated decoders is an important area of future research (Methods). Taking into account the decoding time overhead, we may eventually need to insert more SE rounds to simplify decoding or wait for decoding completion [53], as is also needed for FT protocols that rely on single-shot quantum error correction [25]. In that case, we still expect a significant practical saving over existing schemes. In light of recent experimental advances  [36], a full compilation and evaluation of the space-time savings in parallel reconfigurable architectures such as neutral atom arrays is an important next step. Finally, it will be interesting to investigate how these results can be combined with recent progress toward constant-space-overhead quantum computation [5, 54, 55, 56, 23, 20, 27, 29] or generalized to transversal non-Clifford gates [57, 30, 58, 59, 60, 61, 62], in order to further reduce the space-time volume of large-scale quantum computation.
Author Contributions
H.Z. formulated the decoding strategy and developed an initial proof sketch through discussions with C.Z., M.C., D.B., C.D., S.-T.W., A.K., and M.D.L.. C.Z., M.C., H.Z., and H.H. performed numerical simulations. H.Z., A.K., C.Z., M.C., and C.D. proved the fault tolerance of the scheme. All authors contributed to writing the manuscript.
Acknowledgements
@@
We acknowledge helpful discussions with G. Baranes, P. Bonilla, E. Campbell, S. Evered, S. Geim, L. Jiang, M. Kalinowski, A. Krishna, S. Li, D. Litinski, T. Manovitz, N. Maskara, Y. Wu, and Q. Xu. We would particularly like to thank C. Pattison for early discussions and suggesting the simulation of the $|{\overline{Y}}\rangle$ state distillation factory, and J. Haah for stimulating discussions and deep insights. We acknowledge financial support from IARPA and the Army Research Office, under the Entangled Logical Qubits program (Cooperative Agreement Number W911NF-23-2-0219), the DARPA ONISQ program (grant number W911NF2010021), the DARPA IMPAQT program (grant number HR0011-23-3-0012), the Center for Ultracold Atoms (a NSF Physics Frontiers Center, PHY-1734011), the National Science Foundation (grant number PHY-2012023 and grant number CCF-2313084), the Army Research Office MURI (grant number W911NF-20-1-0082), DOE/LBNL (grant number DE-AC02-05CH11231). M.C. acknowledges support from Department of Energy Computational Science Graduate Fellowship under award number DE-SC0020347. D.B. acknowledges support from the NSF Graduate Research Fellowship Program (grant DGE1745303) and The Fannie and John Hertz Foundation. This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.

### v2

@@ Abstract
Next, we simulate a state distillation factory, which involves the use of noisy state injection and is a key algorithmic subroutine for improving the quality of resource states. We focus on distillation of the $|{\overline{S}}\rangle=\overline{S}|{\overline{+}}\rangle$ state using the [[7,1,3]] Steane code (Fig. 4(a)), which can be efficiently simulated classically and allows the easy implementation of $\overline{S}$ gates in the surface code. Since this circuit has a similar structure to the practically relevant $|{\overline{T}}\rangle$ magic state distillation factories (Extended Data Fig. 4), and the state injection procedure is agnostic to the particular state that is injected, we expect the results to readily generalize to $|{\overline{T}}\rangle$ magic state factories. We fix the error rate of the circuit to $p=0.1\%$, and vary the input infidelity $P_{\mathrm{in}}$ in Fig. 4(c). We post-select on all $X$ logical stabilizers of the Steane code being correct to achieve cubic suppression of logical error rates (Methods). Examining the output $|{\overline{S}}\rangle$ of a one-level factory, we find that as the code distance is increased, the output logical error rate $P_{\mathrm{out}}$ approaches the fidelity expected for ideal Clifford logical gates in the factory $P_{\mathrm{out}}=7P_{\mathrm{in}}^{3}+O(P_{\mathrm{in}}^{4})$ (see Methods for the full expression), across the explored fidelity regime.
@@
Finally, we simulate the logical error rate for a two-level $|{\overline{S}}\rangle$ state distillation factory, involving a total of 113 logical qubits, where the output $|{\overline{S}}\rangle$ states of a $d_{1}=5$ factory is fed into a second factory with $d_{2}=9$, with the distance chosen such that the logical error is dominated by the input state infidelity $P_{\mathrm{in}}=10\%$. As shown in Fig. 4(d), the logical error rates at each level of the distillation procedure are consistent with that expected based on the ideal factory formula. We note that when performing the multi-stage distillation procedure at larger code distances, additional SE rounds may be needed to account for correlated logical errors arising from newly-initialized stabilizers in later stages (SI).
@@
Transversal operations and correlated decoding were recently found to be highly effective in experiments with reconfigurable neutral atom arrays [42]. The principles of transversal algorithmic fault-tolerance described here are the core underlying mechanisms of these observations, such as correlated decoding of a logical Bell state [42]. While recent work has provided strong evidence that this time cost reduction is possible for circuits consisting purely of Clifford gates and Pauli basis inputs [15], up to now it has generally been believed that this conclusion does not hold when performing universal quantum computation [31, 36, 37], which crucially relies on the use of magic states and feed-forward operations. The present work not only demonstrates that this time cost reduction is broadly applicable to universal quantum computing, but also provides a theoretical foundation for it through mathematical fault tolerance proofs.
@@
Although our analysis focused on the use of an MLE decoder, which is inefficient in general, our numerical simulations suggest that algorithms with polynomial runtime can still achieve a competitive threshold [45]. The development of improved, parallel correlated decoders and designing compilations that keep the decoding volume well-controlled is an important area of future research (see Methods for discussion). Indeed, recent analysis [50, 51, 58] shows that surface code transversal algorithms can be decoded using a fast matching decoder, without substantial increase in complexity compared to conventional decoding of individual logical qubits. Taking into account the decoding time overhead, asymptotically, we may eventually need to insert more SE rounds (e.g. $\Theta(\log d)$ rounds for certain fast decoders [59, 60]) to simplify decoding or wait for decoding completion [61], as is also needed for FT protocols that rely on single-shot quantum error correction [26]. In light of recent experimental advances  [42], a full compilation and evaluation of the space-time savings in parallel reconfigurable architectures such as neutral atom arrays is an important next step. Finally, it will be interesting to investigate how these results can be combined with recent progress toward constant-space-overhead quantum computation [5, 62, 63, 64, 24, 21, 28, 30] or generalized to transversal non-Clifford gates [65, 31, 66, 67, 68, 69, 70], in order to further reduce the space-time volume of large-scale quantum computation.
@@
Acknowledgements: We acknowledge helpful discussions with G. Baranes, P. Bonilla, E. Campbell, S. Evered, S. Geim, L. Jiang, M. Kalinowski, A. Krishna, S. Li, D. Litinski, T. Manovitz, Y. Wu, and Q. Xu. We would particularly like to thank C. Pattison for early discussions and suggesting the simulation of the $|{\overline{S}}\rangle$ state distillation factory, and J. Haah for stimulating discussions and deep insights. We would also like to thank M. Beverland, K. Brown, U. Vazirani and the reviewers for suggestions on presentation. We acknowledge financial support from IARPA and the Army Research Office, under the Entangled Logical Qubits program (Cooperative Agreement Number W911NF-23-2-0219), the DARPA ONISQ program (W911NF2010021), the DARPA IMPAQT program (HR0011-23-3-0012), the DARPA MeasQuIT program (HR0011-24-9-0359), the Center for Ultracold Atoms (a NSF Physics Frontiers Center, PHY-1734011), the National Science Foundation (grant number PHY-2012023 and grant number CCF-2313084), the Army Research Office MURI (grant number W911NF-20-1-0082), DOE/LBNL (grant number DE-AC02-05CH11231), the Wellcome Leap Quantum for Bio program. M.C. acknowledges support from Department of Energy Computational Science Graduate Fellowship under award number DE-SC0020347. D.B. acknowledges support from the NSF Graduate Research Fellowship Program (grant DGE1745303) and The Fannie and John Hertz Foundation. This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Author Contributions: H.Z. formulated the decoding strategy and developed an initial proof sketch through discussions with C.Z., M.C., D.B., C.D., S.-T.W., A.K., and M.D.L.. C.Z., M.C., H.Z., and H.H. performed numerical simulations. H.Z., A.K., C.Z., M.C., N.M., and C.D. proved the fault tolerance of the scheme. All authors contributed to writing the manuscript.
Data Availability: The data that supports the findings of this study are available at https://zenodo.org/records/16552626.
Correspondence and requests for materials should be addressed to H.Z. and M.D.L.